# Urban Region Pre-training and Prompting: A Graph-based Approach

### v1

@@ 6.3. RQ3: Manually-designed Prompt Study
6.4. RQ4: Task-learnable Prompt Study
6.4.1. Effectiveness of Task-learnable Prompt.
We utilize GURP pre-trained embedding with the simple prefix prompt (proposed by HREP(Zhou et al., 2023a)) and task-learnable prompt respectively for comparative analysis. Both of these prompt methods are data-driven and seek to improve the region representation by learning data for specific downstream tasks. We conduct experiments on crime prediction task in NYC and crash prediction task in CHI, respectively, and the results are shown in Table 3.
Our task-learnable prompt outperforms pre-trained embeddings across both tasks, with improvements of 8.83% of $R^{2}$ in crime prediction and 149.66% of $R^{2}$ in crash prediction, indicating that the prompt is highly effective in providing tailored guidelines for different downstream tasks. However, in crime prediction task, the prefix prompt have an negative effect, resulting in a performance decrease of 1.16% of $R^{2}$. This demonstrates that our task-learnable prompt is more stable and effective compared to prefix prompt.
Table 3. Effectiveness of Task-learnable Prompt.
@@ 6.4. RQ4: Task-learnable Prompt Study
6.4.2. Visualization of Learned Prompt Graph.
Compared to prefix prompt, our task-learnable prompt is able to generate a series of prompt graphs during training, which enhances the clarity of our prompt method. We visualize prompt graphs that reveal key structural features in the urban region graph of interest in the task. In Figure 8, we show some prompt graphs for crime prediction tasks in CHI. (Additional prompt graphs can be found in Appendix A.8.) These results show that the prompt graphs generally exhibits circular structural features, implying that the model have identified and emphasized interactions between urban entities.
Figure 8. Visualization of learned prompt graphs.
6.5. RQ5: Ablation Studies
Finally, we conduct the ablation studies in NYC dataset to investigate the effectiveness of the individual components of GURPP. Variants of GURPP are designed as follows: GURPP/P. We use only the pre-training process to learn urban region general representation.GURPP/F,  GURPP/S,  GURPP/I. We remove the flow view, spatial proximity view, and imagery view from GURPP, respectively.  GURPP/K. We replace the knowledge graph embedding initialization with random initialization.
@@ 6.5. RQ5: Ablation Studies
Figure 9 presents the ablation results of our model and its variants. Spatial proximity view and knowledge graph embedding method significantly enhance the model’s performance, lifting it by 151.64% and 156.67% of $R^{2}$, respectively. These components effectively capture the heterogeneous and generic patterns of interactions among entities by elucidating the spatial relationships and structured knowledge of the urban region. The flow view and the imagery view can also greatly enhance the performance of the model, increasing it by 29.04% and 21.11%, respectively. This is because they capture dynamic flow information and visual features of urban regions, respectively, providing the model with a multi-dimensional context that enhances its ability to fully understand and predict urban dynamics. Prompt learning, with an improvement of 3.12%, effectively elevates model performance by introducing guiding prompts that allow the model to delve deeper into understanding and utilizing task-relevant knowledge.
Figure 9. Results of ablation studies.
@@ 7. Related Work
Extensive studies have been conducted to learn urban region representation from different perspectives. We combine the existing work categories in Table 4.
Table 4. Cats.of Related Works on Urban Region Embedding
Methods Task-specific learning or fine-tuning Prompting
Non-graph (Yang et al., 2022; Zhou et al., 2023b; Li et al., 2023; Zhang et al., 2023; Li et al., 2024; Yao et al., 2018; Jenkins et al., 2019) (Yuan et al., 2024; Yan et al., 2024; Zhou et al., 2023a)
Graph (Xu et al., 2023; Zhang et al., 2024; Bai et al., 2019; Han et al., 2020; Zhang et al., 2021; Wu et al., 2022; Liu et al., 2023a; Ning et al., 2024) GURPP(Ours)
@@
Task-specific learning or fine-tuning. Urban region representation learning aims to extract complex patterns and relationships within urban regions and embed them into a representation space. Existing research can be broadly classified into two categories: task-specific supervised region representation learning (Bai et al., 2019; Han et al., 2020; Yang et al., 2022; Zhou et al., 2023b) and general unsupervised region representation learning (Li et al., 2023; Zhang et al., 2023; Li et al., 2024; Yao et al., 2018; Wu et al., 2022; Jenkins et al., 2019; Zhang et al., 2021). These works utilize feature aggregation models to generate representation from predefined region statistics with limitations in capturing detailed spatial layout and semantics. Moreover, using predefined statistical features makes it challenging to further adapt the representation to the specificities of various downstream tasks. We address the limitations by leveraging a urban region graph and subgraph-centric region embeddings, which learn the generic knowledge of urban spatial layout and generate flexible region representation tailored to different tasks.
@@
Pre-training and Prompting. Recent years have witnessed the significant success of pre-training and prompting paradigm in fields of NLP (Liu et al., 2023b; Brown et al., 2020; Roberts et al., 2020; Gao et al., 2021), CV (Yoo et al., 2023; Radford et al., 2021; Zhou et al., 2022b, a) and graph learning (Sun et al., 2023; Yu et al., 2024b; Ma et al., 2024; Yu et al., 2024a). Many works in urban computing have also made attempts in this regard (Xu et al., 2023; Zhang et al., 2024; Balsebre et al., 2023; Yuan et al., 2024; Yan et al., 2024). For instance, Yuan et al. (Yuan et al., 2024) design an urban spatio-temporal prediction pre-train and prompt model. Yan et al. (Yan et al., 2024) introduce large language models (LLMs) to describe urban imagery features and unify different prediction tasks through textual prompts. However, the urban tasks explored in these work differ from ours, with a distinct focus. Our work can also benefit from the integration of LLMs to enhance fine-grained characterization capabilities. As a similar work, HREP (Zhou et al., 2023a) introduces prefix prompt to region embedding, but the pre-trained model fails to adequately capture the fine-grained features of regions. We propose both manually-designed and task-learnable prompts built on graph-based pre-train model to address this issue.
@@ 8. Conclusion and Future Works
In this paper, we introduce the urban region graph and propose a graph-based pre-training and prompting framework for urban region representation learning, balancing urban generality and task specificity. We develop a subgraph-centric pre-training model to capture general urban knowledge and two prompt methods to incorporate task-specific insights. Extensive experiments demonstrate the effectiveness of our approach. Future work may explore: 1) Integrating textual data and LLMs to uncover more universal urban knowledge; 2) Developing arbitrary region representations for flexible urban modeling; 3) Assessing the transferability of urban representations across cities to enhance general applicability.
@@ Appendix A Appendix
A.1. Appendix of urban region graph construction
Here we provide more details for the urban region graph construction. Specifically, we incorporate brand information for POI entities, such as Starbucks, to enhance the representation of POI functionalities. Altogether, we define eight node types shared across regions and cities: region, POI, POI category, brand, road, road category, junction, and junction category, comprising the node type set $T_{V}$. We also define six edge types, forming the edge type set $T_{E}$. These include the NearBy relationship, which connects region entities based on their geographical adjacency; the Contains relationship, which is based on spatial containment; the BrandOf relationship, which links brands to POI entities; and the CateOf, JCateOf, and RCateOf relationships, which are based on the category information of the entities.
A.2. Appendix of region imagery encoding
We supplement more details for the urban imagery encoding. Specifically, we employ ResNet (He et al., 2016) to extract visual features for each region image, such that $\vec{a}_{k}=\text{ResNet}(a_{k})$ and derive the comprehensive visual representation for the region by averaging the features of multiple images within the region. Then, a linear projection layer is utilized to project the visual feature into the representation space of region attribute, such that
$\vec{h}_{r}^{img}=\text{Linear}(\frac{1}{m}\sum_{k=1}^{m}\vec{a}_{k}).$
The contrastive loss $\mathcal{L}_{img}$ is then used to align the visual representation with the subgraph-centric region representation, ensuring a cohesive integration of visual and attribute information.
@@ A.2. Appendix of region imagery encoding
$+\log\left(\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{r}^{img}}{\tau}\right)+\sum_{j=1}^{B-1}\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{j}}{\tau}\right)\right)\Big{]}$
Here, $B$ represents batch size and $\tau$ is the temperature parameter.
A.3. Appendix of human mobility encoding
Here we give a detailed description for human mobility encoding. To learn the embeddings, we reconstruct the departure and arrival flow distribution. Let $m_{ij}$ represent the number of trips from region $r_{i}$ to region $r_{j}$. The source and destination urban flow distributions are computed as $P^{{src}}(r_{j}|r_{i})=\frac{m_{ij}}{\sum_{k=1}^{N}m_{ik}}$ and $P^{{dst}}(r_{i}|r_{j})=\frac{m_{ij}}{\sum_{k=1}^{N}m_{kj}}$, respectively. We then reconstruct the source and destination distributions $\hat{P}^{{src}}(r_{j}|r_{i})$ and $\hat{P}^{{dst}}(r_{i}|r_{j})$ using the encoded inflow and outflow features:
$\hat{P}^{{src}}(r_{j}|r_{i})=\frac{\exp((\vec{h}_{i}^{{src}})^{\top}\vec{h}_{j}^{{dst}})}{\sum_{k}\exp((\vec{h}_{i}^{{src}})^{\top}\vec{h}_{k}^{{dst}})}$
$\hat{P}^{{dst}}(r_{i}|r_{j})=\frac{\exp((\vec{h}_{j}^{{dst}})^{\top}\vec{h}_{i}^{{src}})}{\sum_{k}\exp((\vec{h}_{j}^{{dst}})^{\top}\vec{h}_{k}^{{src}})}$
@@ A.3. Appendix of human mobility encoding
$\mathcal{L}_{flow}=\sum_{(r_{i},r_{j})\in\mathcal{R}\times\mathcal{R}}$ $-P^{src}(r_{j}|r_{i})\log\hat{P}^{src}(r_{j}|r_{i})$
$-P^{dst}(r_{i}|r_{j})\log\hat{P}^{dst}(r_{i}|r_{j}),$
which helps in accurately learning the inflow and outflow embeddings by comparing the true distributions with the reconstructed distributions.
A.4. Appendix of attribute graph kernel function
We introduce how attribute graph kernel function $K_{p}(\cdot,\cdot)$ works. We first define the graph kernel function $K(\cdot,\cdot)$. The graph kernel function is a $P$-step random walk kernel (Nikolentzos and Vazirgiannis, 2020) to measure the similarity between graphs, which counts the number of common walks in both graphs within $P$ steps with its calculation as follows. Given $G_{1}=(V_{1},E_{1}),~{}G_{2}=(V_{2},E_{2})$, where $V$ and $E$ are the set of nodes and edges, respectively, we initially define the direct product graph $G_{\times}$ following the framework used in existing research (Nikolentzos and Vazirgiannis, 2020; Feng et al., 2022):
(1) $G_{\times}=G_{1}\times G_{2}=(V_{\times},E_{\times})$
where $V_{\times}=\{(v_{1},v_{2}):v_{1}\in V_{1},v_{2}\in V_{2}\}$, $E_{\times}=\{\{(v_{1},v_{2}),(u_{1},u_{2})\}:\{v_{1},u_{1}\}\in E_{1}\land\{v_{2},u_{2}\}\in E_{2}\}$. Following the theorem (Vishwanathan et al., 2010) that a single random walk on the direct product graph $G_{\times}$ is equivalent to a random walk on the two graphs $G_{1},~{}G_{2}$ at the same time, the $P$-step random walk kernel is computed in the following way:
(2) $K(G_{1},G_{2})=\sum_{p=0}^{P}K_{p}(G_{1},G_{2})=\sum_{p=0}^{P}\lambda_{p}\sum_{i,j=1}^{|V_{\times}|}[A_{\times}^{p}]_{ij}$

### v4

@@ 5.3. Effectivness of Prompting Methods
Results in Table 2 show that focusing on factors such as POI categories, roads, and junctions, closely related to criminal behavior, can improve crime prediction performance. Meanwhile, removing POI brand information shows a significant decrease in the performance of check-in prediction. This demonstrates that our manually-designed prompting method can improve the prediction performance of downstream tasks with explicit knowledge injection.
5.3.2. Task-learnable Prompt Study.
We utilize GURP pre-trained embedding with the simple prefix prompt (proposed by HREP(Zhou et al., 2023a) denoted as GURP+Pre), which directly concatenates hidden vector before pretrained embedding without modeling task-adaptation, and our task-learnable prompt for comparative analysis. Both of these prompt methods are data-driven and seek to improve the region representation for specific downstream tasks. We conduct experiments on crime prediction task in NYC and crash prediction task in CHI, respectively, and the results are shown in Table 3.
Results show that our task-learnable prompt outperforms pre-trained embeddings across both tasks, with improvements of 2.16% ($R^{2}$) in crime prediction and 17.20% ($R^{2}$) in crash prediction, indicating that the prompt is highly effective in providing tailored guidelines for different downstream tasks. In addition, $\text{GURPP}_{T}$ outperforms the prefix prompt in both tasks. This is because the prefix prompt only utilizes the pre-trained region representations for prompting, while our graph-based prompting method also incorporates the original region subgraphs to prompt, better capturing the correlations between task and region attributes.
Table 3. Performance of task-learnable prompt $\text{GURPP}_{T}$.
@@
5.3.3. Comparison of Two Prompting Methods.
We compare the performance of two proposed prompting methods across different cities and tasks to evaluate their effectiveness in various scenarios. In addition to the two prompting patterns P1 and P2, we also define two additional patterns: P3 (randomly samples POI with a weight of 0.9 and ignores road and junction category) and P4 (randomly samples POI with a weight of 0.7 and ignores road and junction), to further validate the effectiveness of different prompting methods. Results show in Table 4.
Table 4. Comparison of two prompting methods.
Method New York City (NYC) Chicago (CHI)
Check-in Carbon Income Crime Carbon Education
RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE
GURP 412.92 0.23 25268 225.24 -0.33 455.43
$\text{GURPP}_{T}$ 388.23 0.17 24745 203.05 0.41 420.62
$\text{GURPP}_{M}$-P1 429.56 -0.13 27839 234.26 -0.35 465.81
$\text{GURPP}_{M}$-P2 397.47 -0.22 26392 234.74 -0.40 458.59
$\text{GURPP}_{M}$-P3 464.90 0.28 28170 256.53 -0.66 486.19
$\text{GURPP}_{M}$-P4 340.25 0.16 22902 238.36 -0.49 502.86
From the results, we observe that when the pre-trained city region data is relatively sparse (e.g., in cases of CHI), task-learnable prompting methods, especially our $\text{GURPP}_{T}$, achieve better enhancement effects. Conversely, when the data is more abundant (e.g., in cases of NYC), manually-designed prompts with explicitly injected knowledge perform better. Moreover, different manually-designed prompts explicitly capture distinct types of knowledge, leading to varying performance across different tasks, which demonstrates the flexibility of our proposed graph-based prompting approach.
5.4. Few-shot & Zero-shot Performance
We compare the performance of $\text{GURPP}_{T}$ and baselines in few-shot and zero-shot scenarios to validate the generalizability of pre-trained representations and superiority of prompting methods.
5.4.1. Few-shot Prediction.
With the pre-trained region representations as input, we randomly sample 10% of the region data as the training set to develop the prediction model, while the remaining 90% is designated as the testing set, to evaluate the performance of $\text{GURPP}_{T}$ with other baselines in few-shot scenarios.
Figure 6(a) shows that $\text{GURPP}_{T}$ outperforms all baselines in the few-shot scenario. This is partly because the pre-trained representations extract generic regional features that exhibit better adaptability in low-resource settings, which, when combined with task-learnable prompts, can better accommodate few-shot prediction. Additionally, our prompting method shows effective transferability compared to HREP.
5.4.2. Zero-shot Prediction.
In this experiment, we first trained the prediction model on NYC crime datasets with NYC pre-trained region embeddings, and then evaluated it on CHI crime datasets with CHI pre-trained region embeddings without any fine-tuning. This setup enables a comparative analysis of the zero-shot transferability between our task-learnable prompt and various baselines. Results show in Figure 6(b).
(a) Few-shot prediction.
(b) Zero-shot prediction.
Figure 6. Few-shot & zero-shot performance.
Compared with baselines without any prompting methods (i.e., HRE, MVURE, and ReCP), models with prompting generally exhibit better zero-shot transferability, suggesting that prompting model can capture task-specific features independent of the city to some extent. Under the same prompting method (prefix prompt), GURP+Pre demonstrates better transferability than HREP, highlighting the superior generalizability of our GURP pre-trained representations. A comparison between GURPP and GURP+Pre reveals that our task-learnable prompting model outperforms the prefix prompt, which can be attributed to the handling of the original region subgraphs within our prompting model.
5.5. Ablation Studies
We perform ablation studies on the NYC dataset to evaluate the effectiveness of each individual component of our proposed model. GURPP includes both pretraining and task-learnable prompting module. GURP contains only the pretraining module. /F, /S, /I, /M denote the removal of the flow view, spatial proximity view, imagery view, and multi-view fusion from GURP, respectively. /K indicates replacing the knowledge graph embedding initialization with random initialization.
@@ 5.5. Ablation Studies
Figure 7 presents the results. Spatial proximity view and knowledge graph embedding method significantly enhance the model’s performance, lifting it by 151.64% and 156.67% of $R^{2}$, respectively. These components capture the heterogeneous and generic patterns of interactions among entities by elucidating the spatial relationships and semantic knowledge of urban regions. The removal of the multi-view fusion module results in learned representations exerting varying effects across different tasks, which indicates that this module contributes to learning more stable representations. The flow view and the imagery view can also greatly enhance the performance of the model, increasing it by 29.04% and 21.11%, respectively. This is because they capture dynamic flow information and visual features of urban regions, respectively, providing the model with a multi-dimensional context that enhances its ability to fully understand and predict urban dynamics. Prompt learning, with an improvement of 3.12%, effectively elevates model performance by introducing guiding prompts that allow the model to delve deeper into understanding and utilizing task-relevant knowledge.
Figure 7. Ablation studies.
5.6. Scalability Analysis
We also validate the scalability of our graph-based method by measuring the pre-training and prompting tuning duration across various scales of regions and region subgraph nodes. Specifically, we use 180 regions in NYC as the base and evaluate the training duration of GURPP over 500 epochs and prompt tuning over 200 epochs for region scale factors of 0.1, 0.2, up to 5 times(1.6 times for pre-training), and evaluate the pretraining duration using 0.05, 0.1, up to 1.0 times the total number of nodes in each region subgraph. For prompt tuning, we conduct evaluations with subgraphs containing 1, 5, up to 120 nodes with the base number 50, which is also the actual experiments settings. The results are averaged over each epoch, shown in Figure 8.
As shown in Figure 8(a), when the region subgraph size is fixed (GURP uses all nodes and prompt tuning uses 50 nodes per subgraph), the training time for pretraining exhibits a superlinear growth with the number of regions. This is mainly due to the flow feature extraction, where the model incorporates dynamics between all region pairs, whereas prompt tuning scales linearly with the number of regions. When the number of regions is fixed, the pretraining duration shows a sublinear relationship with the number of nodes in the regional subgraph, ensuring the scalability of GURP. Due to the random walk kernel, the training time for the task-learnable prompt scales superlinearly with the number of nodes. We set the number of nodes to 50, maintaining acceptable training efficiency.
(a) Pre-training Time.
(b) Prompt Tuning Time.
Figure 8. Impact of data scale on training time.
5.7. Case Study
To validate the superiority of GURPP, we conduct a case study to predict the top 10 regions with the highest crime rates in NYC. The study compares our manually designed prompt, $\text{GURPP}_{M}$-P4, against large the language models (LLMs), including Deepseek R1 (www.deepseek.com) and Gemini 2 Flash (deepmind. google/technologies/gemini/flash/). Detailed descriptions are in Appendix A.5.4.
As shown in the Figure 9, GURPP is able to predict crime regions (the regions in the upper half of the figure) but LLMs fail. This indicates that, compared to language prompts, which directly describe the task to induce relevant knowledge in large language models, GURPP explicitly incorporates task-specific knowledge into the representation using graphs, leading to better performance.
Figure 9. Comparison with language prompt. Shaded regions represent potential crime regions.
@@ 6. Related Work
Urban Region Embedding. Urban region representation learning aims to extract complex patterns and relationships within urban regions and embed them into a representation space. Existing research can be broadly classified into two categories: task-specific supervised region representation learning (Bai et al., 2019; Yang et al., 2022; Zhou et al., 2023b) and general unsupervised region representation learning (Li et al., 2023; Zhang et al., 2023; Li et al., 2024a; Yao et al., 2018; Wu et al., 2022; Jenkins et al., 2019; Zhang et al., 2021; Sun et al., 2024a; Xu and Zhou, 2024). These works, focusing more on exploring multi-view fusion methods to generate representation based on predefined region statistics features with limitations in capturing detailed spatial layout and semantics. Moreover, the lack of fine-grained characterization of region features makes it challenging to further adapt the representation to the specificities of downstream tasks. We address the limitations by using an urban region graph and subgraph-centric region embeddings. Although Zhao et al. (Zhao et al., 2023) employ a graph-based approach to model urban regions at a fine-grained level, their framework is not specifically designed for urban-region pre-training and prompting tasks.
@@
Pre-training and Prompting. The Pre-training and prompting paradigms have achieved significant success in NLP (Liu et al., 2023c; Roberts et al., 2020), CV (Yoo et al., 2023; Zhou et al., 2022b, a), and graph learning (Sun et al., 2023; Yu et al., 2024c; Ma et al., 2024; Yu et al., 2024a; Liu et al., 2023b; Yang et al., 2024; Tang et al., 2024). Many works in urban computing have also made attempts in this regard (Xu et al., 2023; Zhang et al., 2024; Balsebre et al., 2023; Yuan et al., 2024; Yan et al., 2024; Xiao et al., 2024; Li et al., 2024c). For instance, Yuan et al. (Yuan et al., 2024) design an urban spatio-temporal prediction pre-train and prompt model. Yan et al. (Yan et al., 2024) introduce LLMs to describe urban imagery features and unify different prediction tasks through textual prompts. Although the urban tasks they explored differ from ours, their work can inspire ours to integrate LLMs to enhance fine-grained characterization capabilities. HREP (Zhou et al., 2023a) introduces prefix prompt to region embedding, but it lacks explicit extraction of task-adaptation. In this way, we propose both manually-designed and task-learnable prompts built on the graph-based pre-train model to address this issue. The categorization of related work is listed in Table 5.
Table 5. Cats.of related works on urban region embedding.
Category Non-Graph-based Graph-based
Pretraining (Yan et al., 2024; Zhang et al., 2023; Li et al., 2023, 2024a; Xiao et al., 2024) (Fu et al., 2019; Zhang et al., 2021; Wu et al., 2022; Zhou et al., 2023a), GURPP (Ours)
Prompting (Yan et al., 2024; Zhou et al., 2023a) GURPP (Ours)
@@ 7. Conclusion and Future Works
In this paper, we introduce the urban region graph and propose a graph-based pre-training and prompting framework for urban region representation learning, balancing urban generality and task specificity. We develop a subgraph-centric pre-training model to capture general urban knowledge and two prompt methods to incorporate task-specific insights. Extensive experiments demonstrate the effectiveness of our approach. Future work may explore: 1) Integrating textual data and LLMs to uncover more universal urban knowledge; 2) Developing arbitrary region representations for flexible urban modeling (Sun et al., 2025); 3) Assessing the transferability of urban representations across cities to enhance general applicability.
@@ Appendix A Appendix
A.1. Urban Region Graph Construction
Here we provide more details for the urban region graph construction, with the schema shown in Figure 10. We define eight node types shared across regions and cities: region, POI, POI category, brand, road, road category, junction, and junction category in total, comprising the node type set $T_{V}$, and six edge types to form the edge type set $T_{E}$. These include the NearBy relationship, which connects region entities based on their geographical adjacency; the Contains relationship, which is based on spatial containment; the BrandOf relationship, which links brands to POI entities; and the CateOf, JCateOf, and RCateOf relationships, which are based on the category information of the entities.
Figure 10. Schema of urban region graph.
A.2. Region Imagery Encoding
We provide details on urban imagery encoding. ResNet (He et al., 2016) is utilized to extract visual features for each region image, represented as $\vec{a}_{k}=\text{ResNet}(a_{k})$. A comprehensive visual representation for the region is obtained by averaging features from multiple images: $\vec{h}_{r}^{img}=\text{Linear}(\frac{1}{m}\sum_{k=1}^{m}\vec{a}_{k}).$ We then adopt the contrastive loss $\mathcal{L}_{img}$ to align the visual representation with the subgraph-centric embedding.
@@ A.2. Region Imagery Encoding
$+\log\left(\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{r}^{img}}{\tau}\right)+\sum_{j=1}^{B-1}\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{j}}{\tau}\right)\right)\Big{]},$
where, $B$ represents batch size and $\tau$ is the temperature parameter.
A.3. Mobility Pattern Encoding
We provide a detailed description for mobility pattern encoding. We first reconstruct the departure and arrival flow distribution. Let $m_{ij}$ represent the number of trips from region $r_{i}$ to region $r_{j}$. The source and destination urban flow distributions are computed as $P^{{sc}}(r_{j}|r_{i})=\frac{m_{ij}}{\sum_{k=1}^{N}m_{ik}}$ and $P^{{dt}}(r_{i}|r_{j})=\frac{m_{ij}}{\sum_{k=1}^{N}m_{kj}}$, respectively. We then reconstruct the source and destination distributions $\hat{P}^{{sc}}(r_{j}|r_{i})$ and $\hat{P}^{{dt}}(r_{i}|r_{j})$ using the encoded inflow and outflow features:
$\hat{P}^{{sc}}(r_{j}|r_{i})=\frac{\exp((\vec{h}_{i}^{{sc}})^{\top}\vec{h}_{j}^{{dt}})}{\sum_{k}\exp((\vec{h}_{i}^{{sc}})^{\top}\vec{h}_{k}^{{dt}})},\hat{P}^{{dt}}(r_{i}|r_{j})=\frac{\exp((\vec{h}_{j}^{{dt}})^{\top}\vec{h}_{i}^{{sc}})}{\sum_{k}\exp((\vec{h}_{j}^{{dt}})^{\top}\vec{h}_{k}^{{sc}})}.$
@@ A.3. Mobility Pattern Encoding
$\mathcal{L}_{flow}=\sum_{(r_{i},r_{j})}-P^{sc}(r_{j}|r_{i})\log\hat{P}^{sc}(r_{j}|r_{i})-P^{dt}(r_{i}|r_{j})\log\hat{P}^{dt}(r_{i}|r_{j}),$
which aids in learning inflow and outflow embeddings by comparing true and reconstructed distributions.
A.4. Attribute Graph Kernel Function
We introduce how attribute graph kernel function $K_{p}(\cdot,\cdot)$ works. We choose a $P$-step random walk kernel (Feng et al., 2022) as graph kernel function $K(\cdot,\cdot)$ to measure the similarity between graphs, which counts the number of common walks in both graphs within $P$ steps. Given $G_{1}=(V_{1},E_{1}),~{}G_{2}=(V_{2},E_{2})$, where $V$ and $E$ are the set of nodes and edges, we initially define the direct product graph $G_{\times}$ (Feng et al., 2022): $G_{\times}=G_{1}\times G_{2}=(V_{\times},E_{\times}),$ where $V_{\times}=\{(v_{1},v_{2}):v_{1}\in V_{1},v_{2}\in V_{2}\}$, $E_{\times}=\{\{(v_{1},v_{2}),(u_{1},u_{2})\}:\{v_{1},u_{1}\}\in E_{1}\land\{v_{2},u_{2}\}\in E_{2}\}$. Following the theorem (Vishwanathan et al., 2010) that a single random walk on the direct product graph $G_{\times}$ is equivalent to random walk on the two graphs $G_{1},~{}G_{2}$ at the same time, the $P$-step random walk kernel is computed by:
$K(G_{1},G_{2})=\sum_{p=0}^{P}K_{p}(G_{1},G_{2})=\sum_{p=0}^{P}\lambda_{p}\sum_{i,j=1}^{|V_{\times}|}[A_{\times}^{p}]_{ij},$